OpenStudy (natsent06):

Rosa's best time in a race is two seconds faster than her second-best time. Her second-best time is 1 1/4 times her best time. What is her second-best time?

OpenStudy (anonymous):

So what i did was attempt to us algebra to solve based of what the problem tells us. So the first equation is: \[x(b) = 2+ x(sb)\] and \[x(b) = 1 \frac{ 1 }{ 4 }x(sb)\]. Where b = best time and sb = second best time

OpenStudy (anonymous):

From there i subtracted the two equations from each other and got\[0=2-\frac{ 1 }{ 4 }x(sb)\] which goes to \[-2 = -1/4x(sb)\] so \[x(sb)=8\]